OpenStudy (anonymous):

what is the value of the y variable in the solution to the following system of equations? 3x+2y=6 2x+3y=-6

OpenStudy (anonymous):

6 -3 3 -6

OpenStudy (anonymous):

@qweqwe123123123123111 do u know the answer?

OpenStudy (anonymous):

I'm playing with it... :-)

OpenStudy (anonymous):

okay :)

OpenStudy (anonymous):

Okay...ya there?

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Okay. So let's take one of your expressions and figure out an xy equivalency: 3x + 2y = 6 first, let's isolate the Y term by subtracting 3x from both sides. Can you do that?

OpenStudy (anonymous):

subtracting it from 3 and 6?

OpenStudy (anonymous):

No, just the entire 3x term. Subtract it from both sides of the "=" sign.

OpenStudy (anonymous):

1y=6?

OpenStudy (anonymous):

No...let me try showing what I mean this way: To subtract 3x from both sides of 3x + 2y = 6 you do this: 3x - 3x + 2y = 6 - 3x

OpenStudy (anonymous):

2y=3x

OpenStudy (anonymous):

Almost... subtracting the 3x from the left side did leave you with 2y. But subtracting it from the right side doesn't get rid of the 6

OpenStudy (anonymous):

6-3x can u help me with that i dont get how the 6 isnt gone

OpenStudy (anonymous):

Okay, we have to work on both sides of the equal sign. We do that because the left side is equal to the right side. So if we have 3x+2y on the left side, and 6 on the right side, then we want to move the 3x from the left side to the ride side so that we can figure out what "Y" is equal to. So 3x+2y-3x is the same as 3x-3x+2y 3x-3x = 0, so all we have left is 2x but we also have to subtract 3x from the right side, too. The right side is "6", so if we subtract 3x from 6, we have 6-3x So we started out with 3x + 2y = 6 And we subtracted 3x from both sides and ended up with 2y = 6 - 3x Do you see how that works?

OpenStudy (anonymous):

yea so 2y=6-3x

OpenStudy (anonymous):

Right! So now we can DIVIDE both sides by 2. We have to do this because then we'll have an equation that says "Y = " whatever it's equal to. This is part of "solving for Y". So can you divide all the elements on both sides by 2?

OpenStudy (anonymous):

y=3?

OpenStudy (anonymous):

No, 2y/2 = y We also need to know what 6/2 is, and what 3x/2 is

OpenStudy (anonymous):

6/2 is 3 and 3x/2 1.5

OpenStudy (anonymous):

Almost. 6/2 is 3 and 3x/2 =1.5x so dividing 2y = 6 - 3x by 2 is: y = 3 - 1.5x And so now we no what Y is equal to, right?

OpenStudy (anonymous):

1.45?

OpenStudy (anonymous):

*And so now we KNOW what Y is equal to, right? Stupid keyboard... :(

OpenStudy (anonymous):

it cant be 1.45 like i think because thats not an option

OpenStudy (anonymous):

No. Right now, all we have is an equivalency, the equation can't be solved numerically. What we know is that Y = "3-1.5x", so we can use that to replace "Y" in the second equation: 2x + 3y = -6 This is where we can figure out exactly what the value of "X" is. We do this by replacing all the "Y"s with "3-1.5x" So now: 2x + 3y = -6 is seen as: 2x + 3(y) = -6 and becomes 2x + 3(3-1.5x) = -6 rewrite 3(3-1.5x) as 3*3 - 3*1.5x = 9-1.5x This gives you: 2x + 9 - 1.5x = -6 Can you see that?

OpenStudy (anonymous):

oooooooo yea so its -6 thank you alot :)

OpenStudy (anonymous):

Well, yes it is. But we didn't solve for it, you just got lucky. To add insult to injury, I made a typo 2x + 9 - 1.5x = -6 should be 2x + 9 - 4.5x = -6 To finish off (since you probably answered already and got it right), taking: 2x + 9 - 4.5x = -6 we want to put the numbers on one side and the variables on the other, so we should subtract 9 from both sides like this: 2x - 4.5x = -6 - 9 Combining terms gives you: -2.5x = -15 Dividing both sides by -2.5 yields: x = 6 This is our first absolute value. Now we can replace all the "X"s with "6": 2x + 3y = -6 becomes 2*6 + 3y = -6 12 + 3y = -6 subtracting 12 from both sides gives 3y = -6 -12 which is: 3y = -18 Dividing both sides by 3 yields: y = -6 So you see that the answer IS -6, but it wasn't actually figured out until this last step.